OpenStudy (anonymous):

25-(3x+7)less than or equal to 2(x+3)+x

OpenStudy (mertsj):

25-3x-7<=2x+6+x 18-3x<=3x+6 12<=6x 2<=x

OpenStudy (anonymous):

25 - 3x - 7 <= 2x + 6 + x 25 - 6 - 7 <= 2x + 3x + x 12 <= 6x x >= 2